OpenStudy (anonymous):

Determine the constant term in the expansion of (x-2/x^2)^9

OpenStudy (anonymous):

\[(\frac{ x-2 }{ x^2 })^{9}=\frac{ 1 }{ x^18 }*(x-2)^9\] since the maximum power in the expansion of (x-2)^9 is x^9 , and x^9/x^18x=x^-9 then we will be having no constant term!